OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 12 and a13 = -60?

OpenStudy (anonymous):

now you need to find difference fist. do you know how? :)

OpenStudy (anonymous):

no please help?

OpenStudy (anonymous):

ok ... the formula is : \[\LARGE a_{n}=a_1+(n-1)d\] in this case you have to use: \[\LARGE a_{13}=12+(13-1)d\] \[\LARGE -60=12+12d\] \[\LARGE -60-12=12d\] \[\LARGE -72=12d\] \[\LARGE d=-\frac{72}{12}\] \[\LARGE d=-6\] so just substitute now: \[\LARGE a_{n}=a_1+(n-1)d\] \[\LARGE a_{32}=12+[(32-1)\cdot (-6)]\] can you do it now ? :)

OpenStudy (anonymous):

is it -174?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\LARGE a_{32}=12+(-6\cdot 31 )\] \[\LARGE a_{32}=12-186\] \[\LARGE a_{32}=-174\] yes that's correct. Well done :) ...And , Glad to help :)